A rod, 1 cm2 x 200 cm and mass 2 kg, is clamped at its center. When vibrating longitudinally it emits its third overtone in unison with a tuning fork making 6000 vps. How much will the rod be elongated if, when clamped at one end, a stretching force of 980 N is applied to the other end?
1
Expert's answer
2020-07-14T08:51:41-0400

Explanations & Calculations

  • Velocity of the longitudinal waves through a dense medium = "\sqrt {\frac{E}{\rho}}"
  • During the 3rd overtone the length of the rod ("l" )= "\frac{7\lambda}{2}" when the rod is clamped at its center. therefore, the frequency of the third overtone

"\qquad\qquad
\begin{aligned}
\small f &= \small \frac{v}{\lambda}= \frac{7v}{2l} = \frac{7}{2l}\sqrt\frac{E}{\rho}
\end{aligned}"

  • "\small \rho = \large \frac{m}{volume } = \large\frac{2kg}{200\times10^{-6}m^3} = \small 10000kgm^{-3}"


  • Since the rod is in resonance with the tuning fork, f =6000
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small 6000 & = \small \frac{7}{2\times 2m}\sqrt{\frac{E}{10000}}\\
\small E &= \small 1.17551\times 10^{11}Nm^{-2}
\end{aligned}"


  • Therefore, the elongation (e) could be calculated to be,

"\qquad\qquad
\begin{aligned}
\small \frac{F}{A} & =\small E \frac{e}{l}\\
\small e & = \small \frac{F l}{AE}\\
&= \small \frac{980N\times2m}{10^{-4}m^2\times (1.17551\times 10^{11}Nm^{-2})}\\
&= \small \bold{1.667\times 10^{-4}m}\\
&= \small \bold{0.1667mm}
\end{aligned}"